OpenStudy (anonymous):

Okay dividing fractions has always confused me. Exactly how do I divide them?

OpenStudy (anonymous):

keep change to x and filp the fraction

OpenStudy (misty1212):

\[\frac{a}{b}\div\frac{c}{d}=\frac{a}{b}\times \frac{d}{c}\]

OpenStudy (anonymous):

But how would I change it?

OpenStudy (anonymous):

go to youtube dividing frcation song it will help and it is fun

OpenStudy (misty1212):

\[\frac{a}{b}\div\frac{c}{d}=\overbrace{\frac{a}{b}}^{keep}\times \overbrace{\frac{d}{c}}^{flip}\]

OpenStudy (anonymous):

Okay so basically u flip the second fraction and keep the first one the same and multiply it like that?

OpenStudy (anonymous):

Maddie how do you medal

OpenStudy (anonymous):

How do I what?

OpenStudy (anonymous):

give a medal